### v2

@@ Appendix D Additional Details on Language Modeling
We perform all experiments in PyTorch using plainLM [^4] https://github.com/Niccolo-Ajroldi/plainLM pretraining codebase and model implementations. The employed model is a decoder-only $124$M transformer architecture, similar to GPT-2 (Radford et al., 2019; Karpathy, 2022), enhanced with Rotational Positional Embedding (Su et al., 2023), RMSNorm (Zhang & Sennrich, 2019), and SwiGLU (Shazeer, 2020). Models are trained for next-token prediction on 5B tokens from FineWebEdu (Penedo et al., 2024), using a batch size of $512$, context length of $1024$ tokens, and GPT-2 tokenizer. To derive a strong baseline algorithm, we perform a grid search over weight decay, gradient clipping and learning rate schedule hyperparameters, exploring both decay-to-zero and $10\%$ decay, as reported in Table 3. Having identified an optimal baseline across these values, we average on top of it using different LAWA and EMA configurations. We then compare these configurations by means of (A) step-to-validation-target and (B) minimal validation loss achieved throughout training, akin to Section 4 and Section 5 respectively.
@@
Top LR 0.0003, 0.001, 0.003, 0.01, $0.030.0003,\,0.001,\,0.003,\,{\color[rgb]{0.18359375,0.29296875,0.484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.18359375,0.29296875,0.484375}\mathbf{0.01}},\,$ 0.03
@@
Table 4: Language Modeling speed-ups and validation performance. WA provides substantial speed-ups, but minimal generalization gains over the last checkpoint. Results report median and standard deviation over $3$ seeds. The missing standard deviation for the step-to-target is an artifact of the evaluation frequency, which we set to $40$ steps, limiting the resolution of the correspondent metric, and resulting in a standard deviation of zero.
@@ Appendix E Averaging Algorithms
In practice, however, it might be beneficial to use online averaging schemes. As demonstrated throughout this work, averaging can operate as a proxy for a shorter learning rate schedule, providing significantly better performance than the currently available model. Since this aspect has interesting implications for large-scale efficient training, it is worth discussing the possible computational overhead introduced by averaging schemes. For the online versions of EMA or LAWA, we offload the averaging buffer to CPU. We implement the online schemes naively in PyTorch, defaulting to blocking communication, but we note that it is possible to asynchronously offload the model weights to the CPU, using non-blocking communication to update an EMA or a moving average of model weights. We acknowledge a first adoption of this approach in DeepSeekAI (2025), and hope that this work encourages a wider usage of it.
Finally, we note that the poor performance of LAWA and EMA in the inaugural AlgoPerf competition (Kasimbeg et al., 2025) stemmed from the API’s lack of support for switching between model parameters and averaging buffers at evaluation time, resulting in inefficient CPU-GPU transfers on bandwidth-limited hardware. Additionally, AlgoPerf does not support asynchronous transfers, further penalizing such submissions. However, as noted earlier, more efficient implementations are possible in real-world applications, where weight averaging can be used to seamlessly materialize better performing models. We report a PyTorch-style implementation of LAWA and EMA in Algorithms 1 and 2.

### v3

@@ Appendix D Additional Details on Language Modeling
We perform all experiments in PyTorch using ajroldi2024plainlm pretraining codebase and model implementations. The employed model is a decoder-only 124124M transformer architecture, similar to GPT-2 (radford2019language; karpathy2022nanogpt), enhanced with Rotational Positional Embedding (su2023roformerRoPE), RMSNorm (zhang2019rmsnorm), and SwiGLU (shazeer2020glu). Models are trained for next-token prediction on 5B tokens from FineWebEdu (penedo2024finewebedu), using a batch size of $512$, context length of $1024$ tokens, and GPT-2 tokenizer. To derive a strong baseline algorithm, we perform a grid search over weight decay, gradient clipping and learning rate schedule hyperparameters, exploring both decay-to-zero and $10\%$ decay, as reported in Table 3. Having identified an optimal baseline across these values, we average on top of it using different LAWA and EMA configurations. We then compare these configurations by means of (A) step-to-validation-target and (B) minimal validation loss achieved throughout training, akin to Section 4 and Section 5 respectively.
@@
Top LR 0.0003, $003,\,0.001,\,0.003,\,{\color[rgb]{0.18359375,0.29296875,0.484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.18359375,0.29296875,0.484375}\mathbf{0.01}},\,0.0$3
@@
Table 4: Language Modeling speed-ups and validation performance. WA provides substantial speed-ups, but minimal generalization gains over the last checkpoint. Results report median and standard deviation over 33 seeds. The missing standard deviation for the step-to-target is an artifact of the evaluation frequency, which we set to $40$ steps, limiting the resolution of the correspondent metric, and resulting in a standard deviation of zero.
@@ Appendix E Averaging Algorithms
In practice, however, it might be beneficial to use online averaging schemes. As demonstrated throughout this work, averaging can operate as a proxy for a shorter learning rate schedule, providing significantly better performance than the currently available model. Since this aspect has interesting implications for large-scale efficient training, it is worth discussing the possible computational overhead introduced by averaging schemes. For the online versions of EMA or LAWA, we offload the averaging buffer to CPU. We implement the online schemes naively in PyTorch, defaulting to blocking communication, but we note that it is possible to asynchronously offload the model weights to the CPU, using non-blocking communication to update an EMA or a moving average of model weights. We acknowledge a first adoption of this approach in DeepSeekV3, and hope that this work encourages a wider usage of it.
Finally, we note that the poor performance of LAWA and EMA in the inaugural AlgoPerf competition (kasimbeg2025accelerating) stemmed from the API’s lack of support for switching between model parameters and averaging buffers at evaluation time, resulting in inefficient CPU-GPU transfers on bandwidth-limited hardware. Additionally, AlgoPerf does not support asynchronous transfers, further penalizing such submissions. However, as noted earlier, more efficient implementations are possible in real-world applications, where weight averaging can be used to seamlessly materialize better performing models. We report a PyTorch-style implementation of LAWA and EMA in Algorithms 1 and 2.